# Joint moments of characteristic polynomials from the orthogonal and unitary symplectic groups

### v1

@@ 4 Connections to integrable systems
Let $N\geq 1$, $k\geq 2$ and $n_{2},\ldots,n_{k}\geq 0$ be integers. Suppose $a>\sum_{q=2}^{k}qn_{q}-1,b>-1$, and denote by $U(\cdot,\cdot;z)$ the confluent hypergeometric function of second kind. Then, if we define
@@
The above proposition provides a Hankel determinant representation for $\mathbb{E}_{N}^{(a,b)}\Big{[}\mathrm{e}^{-\sum_{j=1}^{N}\frac{t_{1}}{\mathsf{x}_{j}^{(N)}}}\Big{]}$. As an application, we give the proof of Proposition 1.13.
Proof of Proposition 1.13.
@@
By the Lewis Carroll identity we thus obtain:
@@
$\mathbb{E}_{N+1}^{(a-2,b)}\Bigg{[}\mathrm{e}^{-\sum_{j=1}^{N+1}\frac{t_{1}}{\mathsf{x}_{j}^{(N)}}}\Bigg{]}\mathbb{E}_{N-1}^{(a+2,b)}\Bigg{[}\mathrm{e}^{-\sum_{j=1}^{N-1}\frac{t_{1}}{\mathsf{x}_{j}^{(N)}}}\Bigg{]}\frac{C_{N+1}^{(a-2,b)}C_{N-1}^{(a+2,b)}}{(C_{N}^{(a,b)})^{2}}$
@@
Inspired by the idea used in the authors’ previous work [8, Section 4], we tried lifting formula (4.1) to a setting involving partial derivatives of a ”large” Hankel determinant, whose entries are constructed as generating functions of $g_{n}(t_{1})$ with respect to $t_{2},\ldots,t_{k}$, evaluated at $(0,\ldots,0)$. That is,
@@
To overcome this difficulty, we begin by studying a truncated version of the Hankel determinant appearing on the right-hand side of equation (4.1). Specifically, we consider
@@
The motivation behind this truncation approach is from the observation that, with appropriately chosen parameters $R_{2},\ldots,R_{k}$, the basic relation (4.1) is preserved. Specifically, for any $R_{q}\geq n_{q}$ $q=2,\ldots,k$,
@@ 4.2 Hankel determinants shifted by partitions
The result is a mere consequence of the following identities for the confluent hypergeometric functions (see, e.g., [41, p.257 formula (5) and p.258 formula (10)])
@@
Define the partitions $\boldsymbol{\lambda}_{2,1}=(2)$ and $\boldsymbol{\lambda}_{2,2}=(1,1)$. Then, we have Then
@@
[…]
@@
$\begin{pmatrix}\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{2,1}}\\ \mathbf{\Phi}_{N,\boldsymbol{\lambda}_{2,2}}\end{pmatrix}=\begin{pmatrix}1/2&1/2\\ 1/2&-1/2\end{pmatrix}\begin{pmatrix}\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{1,1},1}\\ \mathbf{\Phi}_{N,\emptyset,2}\end{pmatrix}$,
where $\boldsymbol{\lambda}_{1,1}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(1)$. Now, making use of Proposition 4.5, while also noting that $\mathbf{\Phi}_{N,\emptyset,1}=\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{1,1}}$, we have
@@ 4.2.3 A recursive relation for $\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{n,j}}$
The recursive identities for $\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{n,j}}$ take a more concise form when represented in terms of the vectors given as above, where certain vectors will be given as linear combinations of other such vectors upon multiplication by appropriate matrices. Explicit expressions for these matrices are given in the Appendix. In what follows, for the sake of brevity, we write $\partial_{m}$ to mean $\frac{\partial}{\partial t_{m}}$ for any $m=1,\ldots,k$.
@@
$-\frac{\mathbf{A}^{(l)}}{t_{1}+2t_{2}}\sum_{m=0}^{k-3}\mathsf{d}_{m+2,m+3}(-1)^{m}(-N^{2}\partial_{1})\mathbf{\Phi}_{N}^{(1)}[l+m;m+2]$
@@
$+\mathbf{A}^{(l)}\frac{1}{t_{1}+2t_{2}}\sum_{m=2}^{k}\frac{t_{m}^{R_{m}+1}}{N^{2R_{m}}}\tilde{\mathcal{R}}_{m}^{(1)}-\mathbf{A}^{(l)}\frac{kt_{k}}{t_{1}+2t_{2}}\sum_{h=2}^{k}(-1)^{k+h}\frac{t_{h}^{R_{h}}}{N^{2R_{h}}}\tilde{\mathcal{R}}_{k,h}^{(2)}$
$-\mathbf{A}^{(l)}\Bigg{(}\frac{1}{t_{1}+2t_{2}}\Big{(}\sum_{m=3}^{k}(-1)^{m-1}\mathsf{d}_{m-1,m}\sum_{h=2}^{m-1}(-1)^{h}\frac{t_{h}^{R_{h}}}{N^{2R_{h}}}\Big{)}\Bigg{)}\tilde{\mathcal{R}}_{m-1,h}^{(2)},$
@@
Applying Lemma 4.4 and [49, Theorem 17], we get that:
@@
where $\mathcal{W}_{0}\in\mathbb{R}^{l}$ is given by $(\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{l-1,1},1},\dots,\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{l-1,l-1},1},0)$, and the vectors $\mathcal{W}_{m}\in\mathbb{R}^{l}$, defined whenever $m=1,2,\ldots,k+1$, are given by:
@@
Next, we make use of (55) for $\mathcal{W}_{0}$ and (53) for $\mathcal{W}_{m}^{*}$; while for the other terms, we apply [49, Theorem 17 and Proposition 20 ], leading to the identities:
@@
whereas for $k\geq 3$, $2\leq q\leq k$ and $m\in\mathbb{N}$, we have that
@@
First, we let $\boldsymbol{\lambda}=\emptyset$, and using (51), we repeatedly plug in the expression therein for $\mathbf{\Phi}_{N,\boldsymbol{\lambda},l+q-2}$ as $l$ ranges from $3$ to $k+1$. Using this in combination with Lemma 4.5, we get that for $q=2,\ldots,k,m\geq 1$,
@@
In the equation above, to avoid further notational complexity, we assume that for $j=m$, we have
$\sum_{l_{1}=1}^{k-2}\sum_{l_{2}=1}^{k-2}\cdots\sum_{l_{m-j}=1}^{k-2}\Big{(}\prod_{n=1}^{m-j}\big{(}(l_{n}+1)t_{l_{n}+1}-(l_{n}+2)t_{l_{n}+2}\big{)}\Big{)}\mathbf{\Phi}_{N,\emptyset,q+l_{1}+\cdots+l_{m-j}+(k-1)j}=\mathbf{\Phi}_{N,\emptyset,q+(k-1)m},$
and for $k=2,$ $j\leq m-1$,
$\sum_{l_{1}=1}^{k-2}\sum_{l_{2}=1}^{k-2}\cdots\sum_{l_{m-j}=1}^{k-2}\prod_{n=1}^{m-j}\big{(}(l_{n}+1)t_{l_{n}+1}-(l_{n}+2)t_{l_{n}+2}\big{)}\mathbf{\Phi}_{N,\emptyset,q+l_{1}+\cdots+l_{m-j}+(k-1)j}=0.$
The second sum involving $\mathsf{Z}_{b}$’s is also treated similarly. Moving on, upon computation, one can easily show that
@@
where $c_{h_{2},h_{3}^{\prime},\ldots,h_{k-1}^{\prime},h_{k}}^{(i,j)}$ was defined in (59). Even though the quantities above are not well defined for $k=2$, the argument goes through almost verbatim, and hence this case is stated separately in the statement of the result. Finally, observe that the terms in the identities above, other than the term where $h_{2}=0,\ldots,h_{q-1}=0$, can be written in the form of the term involving $V_{i}(t_{1},\ldots,t_{k})$ in the statement of the result. Using these in combination, the desired result now follows. ∎
@@
where $\boldsymbol{\lambda}_{l,q}=(\lambda_{1}^{(l,q)},\ldots,\lambda_{N}^{(l,q)})$ is defined as in Definition (4.2), and
@@
$\mathcal{G}_{N}^{(0,\ldots,0,n_{k})}(t_{1})=\frac{\mathsf{F}_{k-2}^{(0,\ldots,0,n_{k}-1)}(t_{1})}{N^{2}t_{1}}+\sum_{i=1}^{n_{k}-1}\frac{k^{i}}{t_{1}^{i+1}N^{2}}\frac{(n_{k}-1)!}{(n_{k}-1-i)!}\sum_{j=0}^{i}\binom{i}{j}\>c_{0,\ldots,0,i-j}^{(i+1,j)}\mathsf{F}_{k-2+(k-1)i-(i-j)}^{(0,\ldots,0,n_{k}-1-i)}.$
@@
Firstly, for any given $n_{2},n_{3},\ldots,n_{k}$ non-negative integers, we choose $R_{i}=n_{i}+1$ for $i=2,\ldots,k$, and then construct $\mathbf{\Phi}_{N,\boldsymbol{\lambda}}$ as in (46). Then, by means of the following relations that are immediate consequences of their definitions,
@@
$\mathsf{F}_{j}^{(q_{2},\ldots,q_{k})}(t_{1})=\frac{\partial^{q_{2}}}{\partial t_{2}^{q_{2}}}\cdots\frac{\partial^{q_{k}}}{\partial t_{k}^{q_{k}}}\mathsf{Z}_{j}(R_{2},\ldots,R_{k};t_{1},\ldots,t_{k})\Big{|}_{t_{2}=\cdots=t_{k}=0}$
and
$\mathcal{G}_{j}^{(q_{2},\ldots,q_{k})}(t_{1})=\frac{\partial^{q_{2}}}{\partial t_{2}^{q_{2}}}\cdots\frac{\partial^{q_{k}}}{\partial t_{k}^{q_{k}}}\mathbf{\Phi}_{N}(R_{2},\ldots,R_{k};t_{1},\ldots,t_{k})\Big{|}_{t_{2}=\cdots=t_{k}=0}.$
@@
for polynomials $P_{l,j,m}^{(a,b)}(N;t_{1})$, $m\geq 1$, and $t_{1}^{-1}P_{l,j,0}^{(a,b)}(N;t_{1})$ in $t_{1}$ whose degrees do not exceed $l+\sum_{q=2}^{k}qn_{q}-m$ and $l+\sum_{q=2}^{k}qn_{q}-1$, respectively, subject to the additional constraint that they are also polynomials in $N,a,b$, whose degrees in $N$ and in $a,b$ do not exceed $2l+\sum_{q=2}^{k}2(q-1)n_{q}$ and $l+\sum_{q=2}^{k}(q-1)n_{q}-1$, respectively. With this setting in mind, the result then follows by the induction scheme from [8], where we make use of Proposition 4.11 and Proposition 4.12. ∎
@@
Fix $h_{1},\ldots,h_{m}\in\mathbb{N}$, and consider the function,
@@
where $\mathfrak{e}_{N,k}$ is an abbreviation for $\mathfrak{e}_{N,k}^{(a,-\frac{1}{2})}$, as defined in (26) with $a\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{k=0}^{m}h_{k}-\frac{1}{2}$. This is initially well-defined only on $h_{0}\in[\frac{1}{2},\infty)$, thanks to the proof of Theorem 1.3. We claim that it is in fact a well-defined rational function in $h_{0}$ on all of $\mathbb{R}_{\geq 0}$. By Theorem 1.9, we have for $a>\sum_{k=1}^{m}kn_{k}-1$,
@@
$\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[({\mathfrak{p}_{1,N}})^{n_{1}}\prod_{k=2}^{m}(\mathfrak{p}_{q,N})^{n_{k}}\right]=$ $\Bigg{(}\frac{1}{t^{\sum_{k=1}^{m}kn_{k}-1}}\sum_{q=0}^{\sum_{k=2}^{m}(k-1)n_{k}}t_{1}^{q-1}P_{q,1}^{(a,-\frac{1}{2})}(N;t)\frac{\mathrm{d}^{q}}{\mathrm{d}t_{1}^{q}}\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\mathrm{e}^{-t\mathfrak{p}_{1,N}}\right]$
$+\frac{1}{t^{\sum_{k=2}^{m}kn_{k}-1}}\sum_{q=0}^{\sum_{k=2}^{m}(k-1)n_{k}}t_{1}^{q-1}P_{q,2}^{(a,-\frac{1}{2})}(N;t)\frac{\mathrm{d}^{q+n_{1}}}{\mathrm{d}t_{1}^{q+n_{1}}}\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\mathrm{e}^{-t\mathfrak{p}_{1,N}}\right]\Bigg{)}\Bigg{|}_{t=0},$ (71)

### v2

@@ 4 Connections to integrable systems
Let $N\geq 1$, $k\geq 2$ and $n_{2},\ldots,n_{k}\geq 0$ be integers. Suppose $a>\sum_{q=2}^{k}qn_{q}-1,b>-1$, and denote by $U(\cdot,\cdot;z)$ the confluent hypergeometric function of the second kind. Then, if we define
@@
The above proposition provides a Hankel determinant representation for $\mathbb{E}_{N}^{(a,b)}\Big[\mathrm{e}^{-\sum_{j=1}^{N}\frac{t_{1}}{\mathsf{x}_{j}^{(N)}}}\Big]$. As an application, we deduce the Toda lattice equation satified by $\mathbb{E}\left[\mathrm{e}^{-t\mathfrak{e}_{1}(a)}\right]$, which is given as follows.
Proposition 4.2.
Let $a\in\mathbb{R}$ with $a>1$. Then for $t\geq 0$,
Proof.
@@
By the Desnanot-Jacobi identity we thus obtain:
@@
$\mathbb{E}_{N+1}^{(a-2,b)}\Bigg[\mathrm{e}^{-\sum_{j=1}^{N+1}\frac{t_{1}}{\mathsf{x}_{j}^{(N)}}}\Bigg]\mathbb{E}_{N-1}^{(a+2,b)}\Bigg[\mathrm{e}^{-\sum_{j=1}^{N-1}\frac{t_{1}}{\mathsf{x}_{j}^{(N)}}}\Bigg]\frac{(C_{N}^{(a,b)})^{2}}{C_{N+1}^{(a-2,b)}C_{N-1}^{(a+2,b)}}$
@@
Inspired by the idea used in the authors’ previous work [8, Section 4], we tried lifting formula (4.1) to a setting involving partial derivatives of a “large” Hankel determinant, whose entries are constructed as generating functions of $g_{n}(t_{1})$ with respect to $t_{2},\ldots,t_{k}$, evaluated at $(0,\ldots,0)$. That is,
@@
To overcome this difficulty, we begin by studying a truncated version of the Hankel determinant appearing on the right-hand side of equation (43). Specifically, we consider
@@
The motivation behind this truncation approach is from the observation that, with appropriately chosen parameters $R_{2},\ldots,R_{k}$, the basic relation (42) is preserved. Specifically, for any $R_{q}\geq n_{q}$ $q=2,\ldots,k$,
@@ 4.2 Hankel determinants shifted by partitions
The result is a mere consequence of the following identities for the confluent hypergeometric functions (see, e.g., [40, p.257 formula (5) and p.258 formula (10)])
@@
We have
@@
where $\boldsymbol{\lambda}_{n,q}$ and $\mathcal{E}_{2,\emptyset}^{(2)}(t_{1},\ldots,t_{k};R_{2})$ are defined in Definition 4.3 and equation (57),respectively.
@@
$\begin{pmatrix}\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{2,1}}\\ \mathbf{\Phi}_{N,\boldsymbol{\lambda}_{2,2}}\end{pmatrix}=\begin{pmatrix}1/2&1/2\\ 1/2&-1/2\end{pmatrix}\begin{pmatrix}\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{1,1},1}\\ \mathbf{\Phi}_{N,\emptyset,2}\end{pmatrix}$.
Now, making use of Proposition 4.6, while also noting that $\mathbf{\Phi}_{N,\emptyset,1}=\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{1,1}}$, we have
@@ 4.2.3 A recursive relation for $\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{n,j}}$
The recursive identities for $\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{n,j}}$ take a more concise form when represented in terms of the vectors given as above, where certain vectors will be given as linear combinations of other such vectors upon multiplication by appropriate matrices. Explicit expressions for these matrices are given in the Appendix. In what follows, for the sake of brevity, we write $\partial_{m}$ to mean $\frac{\partial}{\partial t_{m}}$.
@@
$+\frac{\mathbf{A}^{(l)}}{t_{1}+2t_{2}}\sum_{m=0}^{k-3}\mathsf{d}_{m+2,m+3}(-1)^{m}(N^{2}\partial_{1})\mathbf{\Phi}_{N}^{(1)}[l+m;m+2]$
@@
$+\mathbf{A}^{(l)}\frac{1}{t_{1}+2t_{2}}\sum_{m=2}^{k}\frac{t_{m}^{R_{m}+1}}{N^{2R_{m}}}\tilde{\mathcal{R}}_{m}^{(1)}-\mathbf{A}^{(l)}\frac{kt_{k}}{t_{1}+2t_{2}}\sum_{h=2}^{k}(-1)^{k+h}\frac{t_{h}^{R_{h}}}{N^{2R_{h}}}\tilde{\mathcal{R}}_{k+1,h}^{(2)}$
$-\mathbf{A}^{(l)}\Bigg(\frac{1}{t_{1}+2t_{2}}\Big(\sum_{m=3}^{k}(-1)^{m-1}\mathsf{d}_{m-1,m}\sum_{h=2}^{m-1}(-1)^{h}\frac{t_{h}^{R_{h}}}{N^{2R_{h}}}\Big)\Bigg)\tilde{\mathcal{R}}_{m,h}^{(2)},$
@@
Applying Lemma 4.5 and [48, Theorem 17],
@@
where $\mathcal{W}_{0}\in\mathbb{R}^{l}$ is given by $(\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{l-1,1},1},\dots,\mathbf{\Phi}_{N,\boldsymbol{\lambda}_{l-1,l-1},1},0)^{T}$, and the vectors $\mathcal{W}_{m}\in\mathbb{R}^{l}$, defined whenever $m=1,2,\ldots,k+1$, are given by:
@@
Next, we make use of (58) for $\mathcal{W}_{0}$ and (56) for $\mathcal{W}_{m}^{*}$; while for the other terms, we apply [48, Theorem 17 and Proposition 20], leading to the identities:
@@
and for $k\geq 3$, $2\leq q\leq k$ and $m\in\mathbb{N}$, we have that
@@
First, we let $\boldsymbol{\lambda}=\emptyset$, and using (54), we repeatedly plug in the expression therein for $\mathbf{\Phi}_{N,\boldsymbol{\lambda},l+q-2}$ as $l$ ranges from $3$ to $k+1$. Using this in combination with Proposition 4.6, we obtain for $q=2,\ldots,k,m\geq 1$,
@@
In the equation above, to avoid further notational complexity, we assume that when $i=2,\ldots,m+1$, for j=i1j=i-1, we have
$\sum_{l_{1}=1}^{k-2}\sum_{l_{2}=1}^{k-2}\cdots\sum_{l_{i-1-j}=1}^{k-2}\Big(\prod_{n=1}^{i-1-j}\big((l_{n}+1)t_{l_{n}+1}-(l_{n}+2)t_{l_{n}+2}\big)\Big)\mathsf{Z}_{q+l_{1}+\cdots+l_{i-1-j}+(k-1)j-2}=\mathsf{Z}_{q+(k-1)(i-1)-2},$
and for $k=2,$ j<i1j<i-1,
$\sum_{l_{1}=1}^{k-2}\sum_{l_{2}=1}^{k-2}\cdots\sum_{l_{i-1-j}=1}^{k-2}\prod_{n=1}^{i-1-j}\big((l_{n}+1)t_{l_{n}+1}-(l_{n}+2)t_{l_{n}+2}\big)\mathsf{Z}_{q+l_{1}+\cdots+l_{i-1-j}+(k-1)j-2}=0.$
The above notation also applies to sums involving $\mathcal{E}_{u,\emptyset,l}^{(1)}$ and $\mathbf{\Phi}_{N,\emptyset,l}$. Moving on, upon computation, one can easily show that
@@
where $c_{h_{2},h_{3}^{\prime},\ldots,h_{k-1}^{\prime},h_{k}}^{(i,j)}$ was defined in (62). Even though the quantities above are not well defined for $k=2$, the argument goes through almost verbatim. Hence this case is stated separately in the statement of the result. Finally, observe that the terms in the identities above, other than the term where $h_{2}=0,\ldots,h_{q-1}=0$, can be written in the form of the term involving $V_{i}(t_{1},\ldots,t_{k})$ in the statement of the result. Using these in combination, the desired result now follows. ∎
@@
where $\boldsymbol{\lambda}_{l,q}=(\lambda_{1}^{(l,q)},\ldots,\lambda_{N}^{(l,q)})$ is defined as in Definition 4.3, and
@@
$\mathcal{G}_{N}^{(0,\ldots,0,n_{k})}(t_{1})=\frac{\mathsf{F}_{k-2}^{(0,\ldots,0,n_{k}-1)}(t_{1})}{N^{2}t_{1}}+\sum_{i=1}^{n_{k}-1}\frac{k^{i}}{t_{1}^{i+1}N^{2}}\frac{(n_{k}-1)!}{(n_{k}-1-i)!}\sum_{j=0}^{i}\binom{i}{j}\>c_{0,\ldots,0,i-j}^{(i+1,j)}\mathsf{F}_{k-2+(k-1)i-(i-j)}^{(0,\ldots,0,n_{k}-1-i)}(t_{1}).$
@@
Firstly, for any given $n_{2},n_{3},\ldots,n_{k}$ non-negative integers, we choose $R_{i}=n_{i}+1$ for $i=2,\ldots,k$, and then construct $\mathbf{\Phi}_{N,\boldsymbol{\lambda}}$ as in (49). Then, by means of the following relations that are consequences of their definitions,
@@
$\mathsf{F}_{j}^{(q_{2},\ldots,q_{k})}(t_{1})=\frac{\partial^{q_{2}}}{\partial t_{2}^{q_{2}}}\cdots\frac{\partial^{q_{k}}}{\partial t_{k}^{q_{k}}}\mathsf{Z}_{j}(R_{2},\ldots,R_{k};t_{1},\ldots,t_{k})\Big|_{t_{2}=\cdots=t_{k}=0}.$
@@
for polynomials $P_{l,j,m}^{(a,b)}(N;t_{1})$, $m\geq 1$, and $t_{1}^{-1}P_{l,j,0}^{(a,b)}(N;t_{1})$ in $t_{1}$ whose degrees do not exceed $l+\sum_{q=2}^{k}qn_{q}-m$ and $l+\sum_{q=2}^{k}qn_{q}-1$, respectively, subject to the additional constraint that they are also polynomials in $N,a,b$, whose degrees in $N$ and in $a,b$ do not exceed $2l+\sum_{q=2}^{k}2(q-1)n_{q}$ and $l+\sum_{q=2}^{k}(q-1)n_{q}-1$, respectively. With this setting in mind, the result then follows by the induction scheme from [8], where we make use of Propositions 4.12 and 4.13. ∎
@@
We first note that when $(h_{1},\ldots,h_{m})=(0,\ldots,0)$, the conclusion in this corollary follows from [46]. We now fix $h_{1},\ldots,h_{m}\in\mathbb{N}\cup\{0\}$ and $(h_{1},\ldots,h_{m})\neq(0,\ldots,0)$, and consider the function,
@@
where $\mathfrak{e}_{N,k}$ is an abbreviation for $\mathfrak{e}_{N,k}^{(a,-\frac{1}{2})}$, as defined in (28) with $a\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{k=0}^{m}h_{k}-\frac{1}{2}$. This is initially well-defined only on $h_{0}\in[\frac{1}{2},\infty)$, thanks to the proof of Theorem 1.3. We claim that it is in fact a well-defined rational function in $h_{0}$ on all of $\mathbb{R}_{\geq 0}$. By Theorem 1.9, we have for $(n_{1},\ldots,n_{m})\neq(0,\ldots,0)$ and $a>\sum_{k=1}^{m}kn_{k}-1$,
@@
$\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[({\mathfrak{p}_{1,N}})^{n_{1}}\prod_{k=2}^{m}(\mathfrak{p}_{q,N})^{n_{k}}\right]=$ $\Bigg(\frac{1}{t^{\sum_{k=1}^{m}kn_{k}-1}}\sum_{q=0}^{\sum_{k=2}^{m}(k-1)n_{k}}t^{q-1}P_{q,1}^{(a,-\frac{1}{2})}(N;t)\frac{\mathrm{d}^{q}}{\mathrm{d}t^{q}}\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\mathrm{e}^{-t\mathfrak{p}_{1,N}}\right]$
$+\frac{1}{t^{\sum_{k=2}^{m}kn_{k}-1}}\sum_{q=0}^{\sum_{k=2}^{m}(k-1)n_{k}}t^{q-1}P_{q,2}^{(a,-\frac{1}{2})}(N;t)\frac{\mathrm{d}^{q+n_{1}}}{\mathrm{d}t^{q+n_{1}}}\mathbb{E}_{N}^{(a,-\frac{1}{2})}\left[\mathrm{e}^{-t\mathfrak{p}_{1,N}}\right]\Bigg)\Bigg|_{t=0},$ (74)